OpenStudy (anonymous):

What is the difference between Initial and Boundary value problems? What is the interval of validity for the following Initial Value Problem. 6(dy/dx) -2y=xy^4; y(0)=-2

OpenStudy (anonymous):

6(dy/dx) -2y=xy^4; y(0)=-2 i have to know the answer of this plsss

OpenStudy (anonymous):

6(dy/dx) -2y=xy^4; y(0)=-2 i have to know the answer of this plsss